OpenStudy (anonymous):

-2x-12 over -2 x − 12 x + 6 −2x + 6 x − 6

OpenStudy (campbell_st):

well it can be written as \[\frac{-2x -12}{-2} = \frac{-2(x + 6)}{-2}\] can you find a common factor to cancel?

OpenStudy (anonymous):

ya the two 2s

OpenStudy (campbell_st):

well it be positive or negative 2

OpenStudy (anonymous):

there wont be any 2s left

OpenStudy (campbell_st):

so whats left...?

OpenStudy (anonymous):

x + 6

OpenStudy (campbell_st):

well done

OpenStudy (anonymous):

ty